The manufacturer’s recommended correct inflation range is 30 psi to 34 psi. Assume the tires’ average psi is on target. If a tire on the car is inspected at random, what is the probability that the tire’s inflation is within the recommended range? (Round your intermediate calculations and final answer to 4 decimal places.)

Probability            
Answer #1

Here no mean and standard deviation

Let μ=32 and σ=3

Since u = 32 and o= 3 we have: ( 30 – 32 XP < < o X-u 34 – 32 P (30 < X <34) =P ( 30 – 32 < X - u < 34 – 32 ) =P 3 Since Z =